# Continual Distillation Learning: Knowledge Distillation in Prompt-based Continual Learning

### v3

@@ 4.2 Limitations of Existing KD Methods in CDL
In prompt-based CL, a new task may select prompt components different from prior tasks, which is an effective strategy to prevent forgetting in continual learning. Because each task can use its own prompts from the prompt pool. However, for knowledge distillation, this mechanism is not effective. For example, in task $A$, a subset of prompt components $\mathcal{P}_{A}$ of the student model is used for learning and knowledge distillation. We hypothesize that these prompt components of the student $\mathcal{P}_{A}$ inherit certain information for the teacher. When task $B$ comes, another subset of prompt components $\mathcal{P}_{B}$ of the student model is selected for learning and knowledge distillation. In the extreme case, if $\mathcal{P}_{A}\bigcap\mathcal{P}_{B}=\emptyset$, the prompts in $\mathcal{P}_{B}$ has no information about the teacher model. Consequently, the student needs to learn from scratch about the teacher model for task $B$.
In knowledge distillation, the goal of the student model is to learn the overall structural knowledge of the teacher model. Distillation should be a global process rather than being confined independently to each task.
@@ 4.3 Knowledge Distillation based on Prompts
To overcome the limitations of previous KD methods in CDL, we introduce a novel knowledge distillation method named Knowledge Distillation based on Prompts (KDP), which is specifically designed for the CDL problem. We introduce a new type of prompts named KD prompts that are globally shared between tasks and are specifically designed for knowledge distillation as illustrated in Fig. 2.
@@
$f_{\text{Pre-T}}(\mathbf{P}^{\text{cl}},\mathbf{h},\mathbf{P}^{\text{kd}})$
$=\text{MSA}(\mathbf{h}_{Q},[\mathbf{P}^{\text{cl}}_{K};\mathbf{h}_{K},\mathbf{P}^{\text{kd}}_{K}],[\mathbf{P}^{\text{cl}}_{V};\mathbf{h}_{V},\mathbf{P}^{\text{kd}}_{V}]).$ (6)
For processing distilled knowledge, we adopt the distillation through attention method in DeiT [31]. In the student model, a distillation token (the KD token in Fig. 2) is inserted at the end of the first layer of the ViT backbone. After being propagated through the ViT backbone, the final output of the distillation embeddings of the KD token are connected to a separate classifier named the KD classifier.
During training, the KD classifier is used to compute the loss function with the teacher classifier using the normal logit distillation loss function. That is, the logits output of the teacher model serves as the “soft labels” to the student. The target is the softened probability distribution of the teacher model, controlled by the temperature factor $\tau$. Therefore, the loss function for the distillation process is:
@@
$+\underbrace{\alpha\mathcal{L}_{\text{KD}}(k_{\phi}^{\mathcal{S}}(f^{\mathcal{S}}_{b}(\mathbf{x};\mathbf{P}^{\text{kd}}_{1}:\mathbf{P}^{\text{kd}}_{n})),g_{\phi}^{\mathcal{T}}(f^{\mathcal{T}}_{b}(\mathbf{x})))}_{\text{student knowledge distillation loss}}$
$+\underbrace{\lambda\mathcal{L}_{\text{pool}}}_{\text{student prompt pool loss}},$ (8)
where $g_{\phi}^{\mathcal{S}}$ and $k_{\phi}^{\mathcal{S}}$ represent the student classifier and KD classifier in the student model, respectively. $g_{\phi}^{\mathcal{T}}$ is the teacher classifier. $f_{b}^{\mathcal{S}}$ and $f_{b}^{\mathcal{T}}$ represent the ViT backbones of the student and teacher models, respectively, which already include the original CL prompts. $\mathbf{P}^{\text{kd}}_{n}$ denotes the KD prompt inserted into the $n$-th block of the ViT backbone. In total, KD prompts are inserted from the first to the $n$-th block. $\mathcal{L}_{\text{pool}}$ is the loss function involved in extracting CL prompts from the prompt pool. Different prompt-based CL methods have different $\mathcal{L}_{\text{pool}}$, as shown in the latter parts of Eq. (2) and Eq. (4). $\alpha$ and $\lambda$ are balancing weights. During testing, the KD classifier and the student classifier are both used for prediction.
@@
5 Experiments
5.1 Datasets
We utilize the CIFAR-100 [18] and ImageNet-R [12] datasets in a class-incremental continual learning setting for our experiments. Following previous works, we divide the ImageNet-R and CIFAR-100 datasets into 10 tasks, where each task contains 10 classes.
5.2 Implementation Details
We experiment with three prompt-based continual learning models, i.e., L2P [35], DualPrompt [34] and CODAPrompt [30]. In L2P, the prompt pool consists of a total of 30 prompt components, and a CL prompt with a length of 20 is inserted only in the first layer of the ViT backbone. In DualPrompt, the prompt pool contains 10 prompt components, each CL prompt has a length of 20, with G-prompts and E-prompts placed in the first two blocks and the third to fifth blocks, respectively. In CODA-Prompt, the prompt pool consists of 100 prompt components. Each CL prompt has a length of 8, and CL prompts are inserted from the first to the fifth layer of the ViT. In our KDP method, unless otherwise specified, all KD prompts have a length of 6 and are inserted from the first block layer to the twelfth layer. Our KDP method by default adopts the DeiT method [31] (including both KD token and the KD classifier) to obtain predictions. This is an optional component that can be added or removed. We evaluate the results of removing the DeiT structure from KDP in the ablation study Sec. 5.5. The parameter $\alpha$ used in Eq. (8) to balance distillation and continual learning is set to 0.5. $\lambda$ for $\mathcal{L}_{\text{pool}}$ is set to 1.
In the experiments, all models are trained only on the data of the current task, and the distillation is based on rehearsal-free models. Our experiments mainly tested two teacher-student knowledge distillation pairs. One is distilling from ViT-Large to ViT-Base [6], and the other one is from ViT-Base to ViT-Small [6].
5.3 Evaluation Metrics
$\text{ACC}=\frac{1}{T}\sum_{i=1}^{T}R_{T,i},$ (9)
$\text{Forgetting}=\frac{1}{T-1}\sum_{i=1}^{T-1}R_{T,i}-R_{i,i}.$ (10)
5.4 Continual Distillation Results
@@
Table 1 summarizes the results of all the aforementioned knowledge distillation methods on the ImageNet-R and CIFAR100 datasets using CODA-Prompt as the continual learner. The complete results for the L2P and DualPrompt methods are provided in the supplementary material. When the teacher model is $\varnothing$, it indicates that there is no knowledge distillation. In the table, DeiT [31] refers to only using the KD token and the KD classifier without incorporating our KD prompts.
From the table, we can see that the logit distillation methods (KD [13], DKD [40]) and feature distillation methods (FitNets [28], ReviewKD [4]) fail to address the limitations discussed in Sec. 4.2. The distilled knowledge is fused into the CL prompt pool, making it difficult to transfer effectively across tasks. As observed in the results, although these four distillation methods can slightly improve the accuracy of the original student model, the improvements are not significant and the forgetting rate remains high. In many cases, the forgetting effect is even worse than that of the original student model. A high forgetting rate indicates that the overall accuracy improvement mainly comes from distillation benefits in the current task, while the influence of distillation on previous tasks is minimal. This further supports our analysis of the limitations in Sec. 4.2.
The DeiT method, using the distillation token and the KD classifier, separates the distilled knowledge from the CL prompt pool. This dual processing approach for CL and KD partially mitigates the interference between task learning and knowledge distillation using prompts. Our KDP method further enhances the separation and cross-task transfer of distilled knowledge through KD-Prompts, ultimately achieving the best overall performance on both datasets as shown in Table 1.
Table 2 presents the distillation results of our KDP method on the L2P, DualPrompt and CODA-Prompt models. This demonstrates the generalization ability of our KDP method in different Prompt-based CL approaches. All three prompt-based CL methods can significantly improve their performance with our KDP method. The combination of CODA-Prompt and our KDP method achieves state-of-the-art (SOTA) performance, particularly benefiting CL with ViT-Small as the backbone. It approaches the performance of larger models while significantly reducing inference computational costs.
@@ 5.5 Ablation Studies
The ablation study is divided into two main aspects. The first aspect explores the number of block layers in which prompts are inserted and the length of each prompt. The second aspect examines the impact of using the KD classifier and the corresponding KD token.
Multiple Layers & KD Prompt Length: For this study, we conducted experiments under the CODA-Prompt and L2P methods. The teacher model backbone is ViT-Base and the student model backbone is ViT-Small. We used the ImageNet-R dataset, where the number of tasks is set to 10. Note that a KD prompt in our model is denoted as $\mathbf{P}^{\text{kd}}\in\mathbb{R}^{L_{p}\times D}$. In the left-hand side of Fig. 3, we show the relationship between the length of the KD prompt, denoted as $L_{p}$, and the accuracy of the KDP model. In this case, the number of inserted layers of the ViT backbone is fixed at 12. In the right-hand side of Fig. 3, we present the relationship between the number of inserted block layers, denoted as $n$, and the accuracy. In this case, $L_{p}=6$.
From the results of the two plots, it can be observed that the length of KD prompts has no strong correlation with the final performance. Therefore, we select $L_{p}=6$, which performs better in CODA-Prompt. As the number of inserted layers $n$ increases, accuracy shows an upward trend in both methods, with a more pronounced effect in L2P. Based on this observation, we set the number of layers to $12$.
@@
Figure 3: Left: The relationship between the length of KD prompt inserted into the model and accuracy. Right: The relationship between the number of layers where KD prompts are inserted and accuracy. Teacher: Vit-Base; Student: ViT-Small; Dataset: ImageNet-R.
@@
KD Classifier: Our method by default adopts the KD classifier structure from DeiT [31] (including the corresponding KD token) to process the soft labels from the teacher. Here, we conducted an experiment to analyze the impact of using the KD classifier.
If neither KD prompts nor the KD classifier is added, the method corresponds to the normal KD [13] in logit distillation. If KD prompts are removed while only adding the KD classifier with the KD token, the method corresponds to DeiT [31]. If both components are incorporated, the method follows the default KDP setup as shown in Fig. 2. Additionally, we design an experiment where only KD prompts are added, in which case the student model has only a single classifier.
As shown in Table 3, the CODA-Prompt method is more heavily relying on the KD classifier for processing. The best performance is achieved when the KD prompts and the KD classifier are used together. However, for the L2P method, the KD classifier does not contribute significantly to improving the distillation. In fact, removing the KD classifier structure leads to an increase in accuracy. However, both CODA-Prompt and L2P depend on KD prompts to achieve better performance, which further validates the effectiveness of the KDP method.
Model KD-Prompts KD-Classifier Avg. Acc ($\uparrow$)
ViT-Small + CODA-Prompt [30] × × $69.91$
ViT-Small + CODA-Prompt [30] × $70.74$
ViT-Small + CODA-Prompt [30] × $70.21$
ViT-Small + CODA-Prompt [30] $\mathbf{71.92}$
ViT-Small + L2P [35] × × $63.97$
ViT-Small + L2P [35] × $64.99$
ViT-Small + L2P [35] × $\mathbf{69.00}$
ViT-Small + L2P [35] $68.18$
Table 3: Ablation study on Split ImageNet-R dataset. Comparison of different models with KD prompts and KD classifier.
@@ Acknowledgments
Supplementary Material
7 Why limiting the scope to prompt-based continual learning?
@@
8 Logit Distillation Details
@@
9 Feature Distillation Details
10 Continual Distillation Results
Tables 8 and 6 present the results of the aforementioned knowledge distillation methods on the ImageNet-R and CIFAR-100 datasets, using L2P [35] and DualPrompt [34] as the continual learners. The results indicate that the KD-Token method consistently optimizes and improves the performance of the original student model across different datasets and various continual learners. Notably, it achieves the highest accuracy in the L2P and DualPrompt models on the CIFAR-10 dataset. Additionally, on the ImageNet-R dataset, the KD-Token method attains the highest accuracy under the base-to-small scheme for both models.

### v4

@@ 4.2 Distillation Information Forgetting
In prompt-based CL, a new task may select prompt components different from prior tasks, which is an effective strategy to prevent forgetting in continual learning. Because each task can use its own prompts from the prompt pool. However, for knowledge distillation, this mechanism is not effective. For example, in task $A$, a subset of prompt components $\mathcal{P}_{A}$ of the student model is used for learning and knowledge distillation. We hypothesize that these prompt components of the student $\mathcal{P}_{A}$ inherit certain information for the teacher. When task $B$ comes, another subset of prompt components $\mathcal{P}_{B}$ of the student model is selected for learning and knowledge distillation. In the extreme case, if $\mathcal{P}_{A}\bigcap\mathcal{P}_{B}=\emptyset$, the prompts in $\mathcal{P}_{B}$ has no information about the teacher model. Consequently, the student needs to learn from scratch about the teacher model for task $B$. In knowledge distillation, the goal of the student model is to learn the overall structural knowledge of the teacher model. Distillation should be a global process rather than being confined independently to each task. As shown in Fig. 3(a), for each new task, the model may re-select different CL prompts. Since the backbone remains frozen, only the CL prompts are updated during training. The distilled knowledge from the previous task is entirely contained in the CL prompts. Re-selecting new CL prompts leads to the loss of teacher information, a phenomenon we refer to as distillation information forgetting.
Since existing distillation methods are not specifically tailored for CDL, the student training process, as illustrated in Fig. 3(a), is prone to the distillation information forgetting problem. A potential solution is to unfreeze part of the ViT backbone (Fig. 3(b)) and allow it to be updated during training. In this way, the backbone updates alongside task progression. However, this approach contradicts the core principle of prompt-based CL, where the backbone is frozen to preserve knowledge from previous tasks. Updating the backbone parameters may further exacerbate forgetting. We have verified this issue in Sec. 5.2. Motivated by the need to address this distillation information forgetting problem, we propose a new solution by introducing KD prompts (Fig. 3(c)).
@@ 4.3 Knowledge Distillation based on Prompts
To overcome the limitations of other KD methods in CDL, we introduce a novel KD method named Knowledge Distillation based on Prompts (KDP), which is specifically designed for the CDL problem. We introduce a new type of prompts named KD prompts that are globally shared between tasks and are specifically designed for knowledge distillation as illustrated in Fig. 2.
@@
$f_{\text{Pre-T}}(\mathbf{P}^{\text{cl}},\mathbf{h},\mathbf{P}^{\text{kd}})=\text{MSA}(\mathbf{h}_{Q},[\mathbf{P}^{\text{cl}}_{K};\mathbf{h}_{K},\mathbf{P}^{\text{kd}}_{K}],[\mathbf{P}^{\text{cl}}_{V};\mathbf{h}_{V},\mathbf{P}^{\text{kd}}_{V}]).$ (4)
For processing distilled knowledge, we adopt the distillation through attention method in DeiT [31]. In the student model, a distillation token (the KD token in Fig. 2) is inserted at the end of the first layer of the ViT backbone. After being propagated through the ViT backbone, the final output of the distillation embeddings of the KD token are connected to a separate classifier named the KD classifier. During training, the KD classifier is used to compute the loss function with the teacher classifier using the normal logit distillation loss function. That is, the logits output of the teacher model serves as the “soft labels” to the student. The target is the softened probability distribution of the teacher model, controlled by the temperature factor $\tau$. Therefore, the loss function for the distillation process is:
@@
$+\underbrace{\alpha\mathcal{L}_{\text{KD}}(k_{\phi}^{\mathcal{S}}(f^{\mathcal{S}}_{b}(\mathbf{x};\mathbf{P}^{\text{kd}}_{1}:\mathbf{P}^{\text{kd}}_{n})),g_{\phi}^{\mathcal{T}}(f^{\mathcal{T}}_{b}(\mathbf{x})))}_{\text{student knowledge distillation loss}}+\underbrace{\lambda\mathcal{L}_{\text{pool}}}_{\text{student prompt pool loss}},$ (6)
where $g_{\phi}^{\mathcal{S}}$ and $k_{\phi}^{\mathcal{S}}$ represent the student classifier and KD classifier in the student model, respectively. $g_{\phi}^{\mathcal{T}}$ is the teacher classifier. $f_{b}^{\mathcal{S}}$ and $f_{b}^{\mathcal{T}}$ represent the ViT backbones of the student and teacher models, respectively, which already include the original CL prompts. $\mathbf{P}^{\text{kd}}_{n}$ denotes the KD prompt inserted into the $n$-th block of ViT. In total, KD prompts are inserted from the first to the $n$-th block. $\mathcal{L}_{\text{pool}}$ is the loss function involved in extracting CL prompts from the prompt pool. Different prompt-based CL methods have different $\mathcal{L}_{\text{pool}}$, as shown in the latter parts of Eq. (1) and Eq. (2). $\alpha$ and $\lambda$ are balancing weights. During testing, the KD classifier and the student classifier are both used for prediction.
5 Experiments
Datasets: We utilize the CIFAR-100 [18] and ImageNet-R [12] datasets in a class-incremental continual learning setting for our experiments. Following previous works, we divide the ImageNet-R and CIFAR-100 datasets into 10 tasks, where each task contains 10 classes.
Implementation Details: We experiment with three prompt-based continual learning models, i.e., L2P [35], DualPrompt [34] and CODAPrompt [30]. In our KDP method, unless otherwise specified, all KD prompts have a length of 6 and are inserted from the first block layer to the twelfth layer. Our KDP method by default adopts both KD prompts and the KD classifier to obtain predictions. We evaluate the results of removing the KD classifier from KDP in the ablation study Sec. 5.2. The parameter $\alpha$ used in Eq. (6) to balance distillation and continual learning is set to 0.5. $\lambda$ for $\mathcal{L}_{\text{pool}}$ is set to 1. In the experiments, all models are trained only on the data of the current task, and the distillation is based on rehearsal-free models. Our experiments mainly tested two teacher-student knowledge distillation pairs. One is distilling from ViT-Large to ViT-Base [6], and the other one is from ViT-Base to ViT-Small [6]. We provide other experimental details in the Appendix B.
Evaluation Metrics: Our experiments use two metrics to evaluate the models: accuracy and forgetting rate [23]. Accuracy refers to the average accuracy of all tasks after completing all 10 tasks. The average forgetting rate, also known as backward transfer, reflects the influence of learning a new task on previously completed tasks. A higher value signifies a more negative impact of the continual learning model. We provide the details in Appendix C.
5.1 Continual Distillation Results
@@
Table 1 summarizes the results of all the aforementioned knowledge distillation methods on the ImageNet-R and CIFAR100 datasets using CODA-Prompt as the continual learner. The complete results for the L2P and DualPrompt methods are provided in the supplementary material. When the teacher model is $\varnothing$, it indicates that there is no knowledge distillation.
As observed in the results, although these logit distillation methods (KD [13], DKD [40]) and feature distillation methods (FitNets [28] can slightly improve the accuracy of the student model, the improvements are not significant and the forgetting rate remains high. In many cases, the forgetting effect is even worse than that of the original student model. A high forgetting rate indicates that the overall accuracy improvement mainly comes from distillation benefits in the current task, while the influence of distillation on previous tasks is minimal. This further supports our analysis of the limitations in Sec. 4.2. The DeiT [31] method, using the distillation token and the KD classifier, separates the distilled knowledge from the CL prompt pool. This dual processing approach for CL and KD partially mitigates the interference between task learning and knowledge distillation using prompts. Our KDP method further enhances the separation and cross-task transfer of distilled knowledge through KD-Prompts, ultimately achieving the best overall performance on both datasets.
Table 2 presents the distillation results of our KDP method on the L2P, DualPrompt and CODA-Prompt models, which demonstrates the generalization ability of our KDP method in different Prompt-based CL approaches (the complete results can be found in Appendix Sec. F). All three prompt-based CL methods can significantly improve their performance with our KDP method. The combination of CODA-Prompt and our KDP method achieves state-of-the-art (SOTA) performance, particularly benefiting CL with ViT-Small as the backbone. It approaches the performance of larger models while significantly reducing inference computational costs.
@@ 5.2 Ablation Studies
We perform ablation studies from multiple perspectives. Unless otherwise specified, all ablation experiments use CODA-Prompt as the baseline method, adopt the ViT-Base to ViT-Small setting for CDL, and are performed on the ImageNet-R dataset.
@@
Figure 4: Left: The relationship between the length of KD prompt and accuracy. Right: The relationship between the number of layers where KD prompts are inserted and accuracy.
Figure 5: The effect of unfreezing the last layer in ViT. After completing all tasks, the test accuracy results for each task are recorded.
Multiple Layers & KD Prompt Length: We conducted experiments under the CODA-Prompt and L2P methods, where the total number of tasks is set to 10. Note that a KD prompt in our model is denoted as $\mathbf{P}^{\text{kd}}\in\mathbb{R}^{L_{p}\times D}$. Fig. 5 (left) shows the effect of KD prompt length $L_{p}$ on accuracy, with the number of inserted layers fixed at 12. Fig. 5 (right) illustrates the impact of the number of inserted layers $n$ on accuracy, with $L_{p}=6$. The results show that $L_{p}$ has limited influence on performance, so we set $L_{p}=6$, which performs well in CODA-Prompt. In contrast, increasing $n$ improves accuracy, especially in L2P, thus we choose $n=12$ for the best performance.
@@
KD Classifier: Our method by default adopts the KD classifier structure (including the corresponding KD token) to process the soft labels from the teacher. Here, we conducted an experiment to analyze the impact of using this structure. If neither KD prompts nor the KD classifier is added, the method corresponds to the normal KD [13] in logit distillation. As shown in Table 3, the best performance is achieved when the KD prompts and the KD classifier are used together.
Unfreezing ViT Backbone: We unfreeze the last layer (block) of the ViT backbone to allow it to participate in training. The experimental results in Fig. 5 support the claim made in Sec. 4.2 that unfreezing the part of ViT backbone does not resolve the Distillation Information Forgetting problem. After completing the training on the final task, we compare the per-task accuracy. The green curve (Unfreeze Last Layer) shows it results in overfitting to the current task and severe forgetting of earlier ones, contradicting the goal of continual learning. The two evaluation metrics, average accuracy dropping to 29.98% and average forgetting reaching 58.21%, indicate that unfreezing the backbone leads to poor overall performance in CDL.
Trainable Prompt Insertion Position: Table 4 investigates the effect of prompt insert position. We add the same number of trainable prompts($L_{p}=6$), originally inserted according to our method, directly into the CL prompt pool. However, this variant is inferior compared to KDP, demonstrating that the improvement is not simply due to the increased number of trainable units. Instead, it shows that our KD prompts contribute to CDL by serving as global, cross-task prompts that effectively mitigate the Distillation Information Forgetting problem.
KD-Prompts KD-Classifier Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
$\varnothing$ $\varnothing$ $69.91$ $7.64$
$\varnothing$ $70.74$ $6.66$
$\varnothing$ $70.21$ $6.98$
$\mathbf{71.92}$ $\mathbf{5.61}$
Table 3: Comparison of the performance of the KDP method under different settings, based on whether KD prompts or the KD classifier are included.
Position of additional prompts Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
CL-Prompt + CODA-Prompt [30] $69.34$ $8.84$
KD-Prompt (Ours) + CODA-Prompt [30] $\mathbf{71.92}$ $\mathbf{5.61}$
CL-Prompt + DualPrompt [34] $65.98$ $6.82$
KD-Prompt (Ours) + DualPrompt [34] $\mathbf{68.77}$ $\mathbf{3.13}$
Table 4: Comparing the insertion positions of the additional prompts. CL-Prompt refers to inserting an equal number of additional trainable prompts into the CL prompt pool.
@@ Acknowledgments
Appendix
Appendix A Why limiting the scope to prompt-based continual learning?
@@
Appendix B Implementation Details
In L2P, the prompt pool consists of a total of 30 prompt components, and a CL prompt with a length of 20 is inserted only in the first layer of the ViT backbone. In DualPrompt, the prompt pool contains 10 prompt components, each CL prompt has a length of 20, with G-prompts and E-prompts placed in the first two blocks and the third to fifth blocks, respectively. In CODA-Prompt, the prompt pool consists of 100 prompt components. Each CL prompt has a length of 8, and CL prompts are inserted from the first to the fifth layer of the ViT. In our KDP method, unless otherwise specified, all KD prompts have a length of 6 and are inserted from the first block layer to the twelfth layer. Our KDP method by default adopts the DeiT method [31] (including both KD token and the KD classifier) to obtain predictions. This is an optional component that can be added or removed. We evaluate the results of removing the DeiT structure from KDP in the ablation study Sec. 5.2. The parameter $\alpha$ used in Eq. (6) to balance distillation and continual learning is set to 0.5. $\lambda$ for $\mathcal{L}_{\text{pool}}$ is set to 1.
Appendix C Evaluation Metrics
Our experiments use two metrics to evaluate the models: accuracy and forgetting rate [23]. Accuracy refers to the average accuracy of all tasks after completing all 10 tasks.
$\text{ACC}=\frac{1}{T}\sum_{i=1}^{T}R_{T,i},$ (7)
$\text{Forgetting}=\frac{1}{T-1}\sum_{i=1}^{T-1}R_{T,i}-R_{i,i}.$ (8)
Appendix D Logit Distillation Details
@@
Appendix E Feature Distillation Details
Appendix F Continual Distillation Results
@@
Tables F and 7 present the results of the aforementioned knowledge distillation methods on the ImageNet-R and CIFAR-100 datasets, using L2P [35] and DualPrompt [34] as the continual learners. The results indicate that the KD-Token method consistently optimizes and improves the performance of the original student model across different datasets and various continual learners. Notably, it achieves the highest accuracy in the L2P and DualPrompt models on the CIFAR-10 dataset. Additionally, on the ImageNet-R dataset, the KD-Token method attains the highest accuracy under the base-to-small scheme for both models.
Appendix G Broader Impacts
This work proposes KD-Prompt (KDP), a new framework for Continual Distillation Learning (CDL), which enables scalable knowledge transfer from large teacher models to compact student models across a sequence of tasks. In many recognition and segmentation scenarios, large vision models achieve strong performance but are impractical for deployment on resource-constrained platforms. Our method addresses this by continually distilling knowledge from the high-capacity teacher into a lightweight student, thus supporting continual learning while maintaining inference efficiency.
As new tasks arrive, the teacher model acts as a continuously evolving base model, and the task-adapted student model obtained through continual distillation is used for practical deployment without catastrophic forgetting. This approach enables deploying compact, efficient models on latency-sensitive platforms such as mobile devices and robots, where low inference cost and fast response time are essential, while also supporting the need for continual learning.
However, this process introduces trade-offs. Training large teacher models for each new task may incur substantial energy and computational costs. Therefore, we emphasize the need for responsible deployment, where the benefits of using smaller student models in deployment environments must be weighed against the training overhead of large-scale teacher models.
In summary, KDP enables continual task adaptation and compression of high-performing models into compact ones suitable for real-world applications, making progress toward sustainable and practical continual learning systems.